OpenStudy (anonymous):

what is the domain of \[1/\sqrt[4]{x ^{2}-2x}\]

myininaya (myininaya):

so you want to solve the following: \[x^2-2x >0\]

myininaya (myininaya):

\[f(x)=x^2-2x=x(x-2) => f=0 \text{ when } x=0,2\] Now I will just test the areas around x=0 and x=2 to see where f>0

OpenStudy (anonymous):

myiniyaya about the other problem...i did as you said but i just cant seem to solve these two [f(a)]^2 2f(a)

OpenStudy (anonymous):

they are very strange

myininaya (myininaya):

---|----|---- 0 2 To test interval (-inf,0) we need to plug in a value of this interval into f Samething for (0,2) and (2,inf)